This problem in extra work but i am confused. If anyone can help thanks.
Solve algebraically, using the five step method. Show all steps.
The area of a rectangle in 525cm^2 Find the dimensions of a rectangle if the ratio of the length to the width is 7:3

Answer by ankor@dixie-net.com(22740)

The area of a rectangle in 525cm^2
Find the dimensions of a rectangle if the ratio of the length to the width is 7:3
:
let x = the multiplier
then
7x = the length
and
3x = the width
:
Area equation (L*W)
7x * 3x = 525
21x^2 = 525
Divide both sides by 21, results
x^2 = 25
x = 5 is the multiplier
then
5*7 = 35 is the length
and
5*3 = 15 is the width
;
Check; 35 * 15 = 525